payton:

Consider the equation 4 plus v equals 8 minus StartFraction 5 Over 3 EndFraction v.v + 4 + StartFraction 5 Over 3 EndFraction v plus 4 equals 8 minus StartFraction 1 Over 3 EndFraction v. v = 8. What is the resulting equation after the first step in the solution? v +4 = 8 – v StartFraction 4 Over 3 EndFraction v plus 4 equals 8.v +4 = 8 4 + v = 8 – 4 plus v equals 8 minus StartFraction 5 Over 3 EndFraction v.v 2v + 4 = 8

TheSmartOne:

You have to simplify it Do you know how to subtract fractions? \(\frac{a}{c} - \frac{b}{c} = \frac{a-b}{c}\)

TheSmartOne:

Oh wait it's addition, similar concept but you just add the numerators ONLY IF the denominator is the same

TheSmartOne:

And in this case when you add it, you can simplify the fraction too

payton:

would it be 5/3v+4=8- 1/3v

TheSmartOne:

So that is us moving it to the other side, but we don't want to do that Why? Because we want to keep the stuff with the variable on one side and to add them together so that way we can get that letter (aka variable) all by itself

TheSmartOne:

What is \(\frac{5}{3}v + \frac{1}{3}v = ?\)

payton:

it should be v=2

TheSmartOne:

are you sure about that? how'd you get that

payton:

i added the 5/3v and the 1/3v

payton:

or is it 2v

TheSmartOne:

mhmm and then you went to v=2

TheSmartOne:

Yes it has to be 2v

TheSmartOne:

Imagine v is just like a cake You have 5/3 of a cake and you get another 1/3 piece of cake Now you have a full 6/3 or 2 cakes So just like that (5/3)v + (1/3)v = (6/3)v = 2v

TheSmartOne:

So now that you got 2v, you should be able to get your answer :)

payton:

2v + 4 = 8

TheSmartOne:

That's your answer!

payton:

i couldnt have done any of this without you

payton:

thanks

TheSmartOne:

Of course!! :)